OpenStudy (anonymous):

Which number is an integer, a rational number, and a real number? A. –14 B. 0.1 C. 2/3 D. 3.14

OpenStudy (tkhunny):

All Real. You don't see any \(i\) or \(j\) in there, do you? All Rational -14 = -14/1 <== Ratio of two integers! 0.1 = 1/10 <== Ratio of two integers! 2/3 <== Ratio of two integers with no alteration! 3.14 = 314/100 <== Ratio of two integers! What do you think of integers? Any candidates?

OpenStudy (anonymous):

Thanks^_^